cathecat88:

Four friends share 1/4 of a gallon of lemonade equally. What fraction of the gallon of lemonade does each friend get?

supie:

Do \(4 \div \frac14\)

snowflake0531:

1/4 divided by 4 people

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Do \(4 \div \frac14\) \(\color{#0cbb34}{\text{End of Quote}}\) wrong

supie:

Oh wait- i got it the wrong way-

supie:

\(\frac14 \div 4\) ?

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(\frac14 \div 4\) ? \(\color{#0cbb34}{\text{End of Quote}}\) Much better

snowflake0531:

LOL btw, @ person who asked the question, do you know hwo to divide fractions/multiply

AZ:

\(\dfrac{a}{b} \div \dfrac{c}{d} = \dfrac{a}{b} \times \dfrac{d}{c}\) remember 4 can be re-written as \(\dfrac{4}{1}\) and lastly \( \dfrac{a}{b} \times \dfrac{d}{c} = \dfrac{ad}{bc}\)

AZ:

so you have to flip the second fraction and then you can multiply the fractions

snowflake0531:

a flawless answer as usual~